Diogo has a utility function𝑈[𝐵, 𝑍] = 𝐴𝐵∝𝑍

𝛽, where A, α and β are constants, B is burritos, and Z is pizzas. If 

the price of burritos, Pb, is N$2 and the price of pizzas, Pz is N$1, and Y is N$100, what is Diogo’s optimal 

bundle?


1
Expert's answer
2021-04-12T13:09:36-0400
"2B+1Z=100"

"\frac {\alpha A B^{\alpha-1}Z}{2}=\frac {AB^{\alpha}}{1}"